OpenStudy (anonymous):

Write the first five terms of the sequence. an = n 2 + 2

Parth (parthkohli):

To find the first term, replace n with 1. Second - n = 2 Third - n = 3 Fourth - n = 4 Fifth - n = 5

OpenStudy (anonymous):

Put n= 1 to 5 each time and get the terms...

Parth (parthkohli):

\( \color{Black}{\Rightarrow a_1 = 1^2 + 2 = 1 + 2 = 3}\) \( \color{Black}{\Rightarrow a_2 = 2^2 + 2 = 4 + 6 = 10}\) etc.